1. You have genotyped an entire population of lab mice for a single locus (let's label it F) that you have noticed have some interesting differences in size at a certain age. Based on your genotyping work, you calculate the average of the FF genotype to be 71 grams, the Ff genotype to be 50 grams and the ff genotype to be 29 grams. What is a in grams for this population (rounded to the nearest 0.1)?
Answer #1

Solution :

Given data is :

Based on genotype work,

Average of the FF genotype to be 71 grams,

average of the FF genotype to be 50 grams,

average of the FF genotype to be 29 grams

Here population size = 71 + 50 + 29 = 150

therefore,

frequency of F allele A says P = [2(71) + 1(50)] / 2(150)

= [ 142 + 50] / 300

= 192 / 300

= 0.64

p = 0.64

we already know that q = 1 - p = 1 - 0.64 = 0.36

f = 0.32